D Question 36 1 pt Caskets are produced in a monopolistic competitive market. One producer, Final Boxes, sells 20 caskets a week at a price of $550 each. Its average total cost is $600. From this information, we know that O this producer is setting marginal revenue equals marginal cost. O this producer is making an economic profit of $500. O new casket firms will want to enter O this producer is losing $1,000 a week. O this producer should increase production. Previous ex e閊曲
DQuestion 39 Firms in a monopolistically competitive industry produce O monopolistic goods only. O only industrial products-and no consumer products O differentiated products. O homogeneous goods and services. O only consumer products-and no industrial products 4 Previous Quiz saved at 9:12pm
D Question 41 1 pts An agreement between Nike and Adidas to raise prices of the track shoes that each company produces by 50 percent is an example of a collusive agreement, and economists generally agree that O this agreement is in the best interest of society because the price of track shoes is significantly above marginal cost. O this agreement is not in the best interest of society because the price of track shoes is significantly above marginal cost. O this agreement is not in the best interest of society because the price of track shoes is significantly below marginal cost O the price of track shoes does not affect societal welfare o this agreement is in the best interest of society because the quantity of track shoes bought and sold is significantly less than the quantity that would be bought and sold in a perfectly competitive market Previous Next e n
media%2F6cf%2F6cfab948-fcd7-4776-807a-e2
There are four Jet Ski rental companies on a small tourist island. The following table shows the quantity of Jet Skis that each firm rents in a typical year and the per-hour price that each firm currently charges for each Jet Ski it rents. If Nautical Jet Skis raised its price to $60.25 and all other firms subsequently increased their prices by $10.00, all firms have Firm Wide Island Waves Nautical Jet Skis Quantity 50 1,200 78 140 Price $50.00 51.25 51.69 50.99 Rapid Ocean Rentals Ocean Rider Emporium en antitrust laws, even if they might not have participated in explicit price collusion. O likely participated in tacit collusion and the government can improve social welfare by banning Jet Ski rentals on the island. O likely participated in explicit price collusion and they all have broken antitrust laws. O likely participated in tacit collusion and they all have broken antitrust laws O likely participated in tacit collusion and they have not broken antitrust laws because tacit collusion is not illegal. a Previous Next
Answer #1

D. This producer is losing $1000 a week.

20 caskets at $550

Total revenue = 20*550= $11000

Average total cost $600

Total cost = 20*600= $12000

Profit = TR-TC= $11000-$12000=-$1000

Or loss= $1000